# Gravitational waves from primordial black hole dominance: The effect of inflaton decay rate

### v1

@@ Abstract
In this work, we explore primordial black holes (PBH) formation scenario during the post-inflationary preheating stage dominated by the inflaton field. We consider, in particular, a model-independent parametrization of the Gaussian peak inflationary power spectrum that leads to amplified inflationary density fluctuations before the end of inflation. These modes can reenter the horizon during preheating and could experience instabilities that trigger the production of PBH. This is estimated with the Khlopov-Polnarev (KP) formalism that takes into account non-spherical effects. We derive an accurate analytical expression for the mass fraction under the KP formalism that fits well with the numerical evaluation. Particularly, we focus on ultra-light PBH of masses $M_{\text{PBH}}<10^{9}g$ and study their evolution and (possible) dominance after the decay of the inflation field into radiation and before the PBH evaporation via Hawking radiation. These considerations alter the previous estimates of induced gravitational waves (GWs) from PBH dominance and open new windows for detecting stochastic GW backgrounds with future detectors.
@@ 1 Introduction
The inflationary phase constitutes the most important phase of the primordial Universe. It solves the conceptual challenges of the classical Big Bang Theory and provides a mechanism for generating the primordial density fluctuations observed in the cosmic microwave background (CMB). These fluctuations act as seeds for many phenomena, including primordial black holes (PBH) Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx ; Carr:1975qj and gravitational waves (GWs) Guzzetti:2016mkm ; Baumann:2009ds ; Maggiore:2007 ; Dodelson:2003 ; Gravitation:1973 ; Franciolini:2021 ; Domenech:2021ztg ; Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 , in posterior eras. The specific nature of each era (matter or radiation-dominated) influences the dynamics of these phenomena, providing insights into early universe physics.
@@
Almost every inflationary model, especially single-field models, exhibits a phase known as preheating Kofman:1994rk ; Kofman:1997yn , just after the end of inflation. During this stage, the universe undergoes a transient matter-dominated period driven by the oscillations of the inflaton field at the bottom of the inflationary potential. PBH are usually considered to form from the collapse of amplified density perturbations generated during inflation as they enter the horizon. Conventionally, PBH formation has been studied in radiation-dominated eras, where the radiation pressure allows the definition of a collapse threshold for the perturbations Carr:2020gox ; Harada:2013epa ; Musco:2012au , and the mass fraction of PBH is often estimated using the Press-Schechter formalism Press:1973iz ; Harada:2013epa . In contrast, during this preheating stage, characterized by negligible pressure, the collapse dynamics differ significantly. Here, non-spherical effects become the dominant factor in stopping the collapse, a scenario well described by the Khlopov-Polnarev (KP) formalism Khlopov:1980mg ; Polnarev:1981 ; Khlopov:1982ef ; Polnarev:1985btg ; Khlopov:2008qy ; Harada:2016mhb . This last, typically ignored in the literature, was originally thought for matter-dominated scenarios. We choose to work under this formalism as we understand that it better describes the collapse dynamics of perturbations in matter-dominated scenarios. We have already explored this in previous works del-Corral:2023apl ; draft-alpha-PBH , but here, we consider two new elements: the decay of the inflation field into radiation and the posterior evaporation of the PBH via Hawking radiation, which we describe in what follows. Further, we will apply these considerations to an extended distribution of perturbations, contrary to the standard monochromatic assumption.
@@
Since PBH form from the collapse of amplified perturbations, a central element in its formation is the appearance of a pronounced peak in the inflationary power spectrum. Various inflationary scenarios can lead to such an amplification of curvature perturbations. For instance, in single-field models, mechanisms such as an inflection point Iacconi:2021ltm ; Dalianis:2018frf ; Garcia-Bellido:2017mdw ; Hertzberg:2017dkh or a step-like feature in the potential Inomata:2021tpx ; Dalianis:2023pur ; Dalianis:2021iig ; Cai:2021zsp can generate the necessary peak. For multi-field, this peak can be produced through non-canonical kinetic terms Aldabergenov:2022rfc ; Braglia:2020eai ; Pi:2021dft ; Wang:2024vfv , the interplay of multiple axion fields Zhou:2020kkf ; Mavromatos:2022yql , or waterfall trajectories in hybrid models Afzal:2024xci ; Spanos:2021hpk ; Braglia:2022phb ; Clesse:2015wea ; Tada:2023fvd . For a comprehensive review of these mechanisms, see Stamou:2024lqf . Although the parametrization of the power spectrum via a Gaussian peak is introduced in an ad hoc manner in our study, adjusting its height and position effectively encapsulates the essential physics of almost every inflationary model mentioned earlier. This is the approach we will follow, remaining agnostic about the precise physical mechanism of the amplification of perturbations during inflation.
@@
Interest in PBH has grown recently due to their potential to answer several questions in cosmology. In particular, they have been proposed as candidates for dark matter, as generators of structure in the universe, or even as seeds for the formation of supermassive black holes in the center of galactic nuclei. See Villanueva-Domingo:2021spv ; Carr:2020gox for a review. However, although observational data tightly constrain their abundance Carr:2020gox ; Cai:2021zsp ; Markov:1984xd ; Coleman:1991ku ; Carr:1994ar ; Zeldovich ; Acharya:2020jbv ; Carr:2009jm ; Josan:2009qn ; Carr:2020xqk , for small masses ($\lesssim 10^{10}$g), these constraints rely on the nature of Dark Matter, which is currently unknown Allahverdi:2020bys . For this reason, the constraints can be relaxed to the point that PBH come to dominate the energy density of the universe, a period called PBH-dominated, which behaves as pressureless matter. If this occurs, their inherent Poissonian fluctuations in density can source a stochastic background of gravitational waves, providing a unique observational window into the early universe.
@@
To eventually recover the standard radiation-dominated era necessary for successful Big Bang Nucleosynthesis (BBN), we consider the inflaton decay into Standard Model particles, thereby reheating the universe. This is achieved by considering the scalar field’s decay into radiation, with decay rate $\Gamma_{\phi}$. However, if PBH dominates before that, then the universe is reheated through Hawking evaporation Hawking:1974rv ; Martin:2019nuw ; Gondolo:2020uqv , and the reheating temperature corresponds to the evaporation temperature of the PBH. Refs. Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2024wao ; Domenech:2020ssp ; Papanikolaou:2022chm considered that the PBH-dominated phase occurs due to the interplay between the different energy densities of PBH (matter), and a background fluid with an arbitrary equation of state $w>0$, ignoring the specific formation mechanism of the PBH. We, however, consider that PBH are produced during an early matter-dominated phase, under the KP formalism, and that the PBH-dominated phase occurs due to the decay of the scalar field into radiation, which modifies the actual constraints on the abundance of PBH. Then, we analyze the production of induced GWs by the Poissonian fluctuations of the PBH fluid. To our knowledge, this is the first time it has been studied in the literature.
@@ 2 Inflation and preheating
where $\mathcal{A}_{s}=2.1\times 10^{-9}$ is the amplitude of the power spectrum at the pivot scale $k_{0}=0.05\text{Mpc}^{-1}$, and $n_{s}\simeq 0.965$ is the spectral index. Motivated by other works Iacconi:2021ltm ; Dalianis:2018frf ; Stamou:2024lqf ; Garcia-Bellido:2017mdw ; Hertzberg:2017dkh ; Inomata:2021tpx ; Dalianis:2023pur ; Dalianis:2021iig ; Cai:2021zsp ; Aldabergenov:2022rfc ; Braglia:2020eai ; Pi:2021dft ; Wang:2024vfv ; Zhou:2020kkf ; Mavromatos:2022yql ; Afzal:2024xci ; Spanos:2021hpk ; Braglia:2022phb ; Clesse:2015wea ; Tada:2023fvd (see the introduction), we introduce a Gaussian peak in the power spectrum, and choose to parametrize it as a log-normal-like distribution, given by
@@
where $\mathcal{A}_{\text{peak}}$ is the amplitude at the peak scale $k_{\text{peak}}$ and the standard deviation of the Gaussian is chosen to be $\sigma=0.2$ to reduce the number of free parameters. Eqn. (2) is useful for this work as it avoids the tails of the Gaussian peak (due to the $\log_{10}$ term), and since it constitutes a two-parameter ($\mathcal{A}_{\text{peak}}$ and $k_{\text{peak}}$) model-independent parametrization that can reflect the physics of the models described in Iacconi:2021ltm ; Dalianis:2018frf ; Garcia-Bellido:2017mdw ; Hertzberg:2017dkh ; Inomata:2021tpx ; Dalianis:2023pur ; Dalianis:2021iig ; Cai:2021zsp ; Aldabergenov:2022rfc ; Braglia:2020eai ; Pi:2021dft ; Wang:2024vfv ; Zhou:2020kkf ; Mavromatos:2022yql ; Afzal:2024xci ; Spanos:2021hpk ; Braglia:2022phb ; Clesse:2015wea ; Tada:2023fvd ; Stamou:2024lqf . The total power spectrum is the sum of the contributions from (1) and (2), that is
@@
During this phase, the perturbations that enter the horizon collapse into PBH, as shown in the next section. However, the quantity we use to compute the abundance of PBH, i.e., the mass fraction $\beta(k)$, is not the power spectrum of curvature perturbations, but the variance of the density perturbations, $\sigma_{k}$. For modes that enter the horizon after inflation, the curvature perturbations are related to the density perturbations $\delta_{\phi,\bm{k}}$ as follows  [^1] In this work, since our focus is only on the scales that are amplified during inflation, we do not consider the subhorizon modes that never exit the horizon during inflation. Although they could contribute to PBH formation, as shown in del-Corral:2023apl ; draft-alpha-PBH , their contribution is only toward very small scales in contrast to what we explore in this work.Martin:2019nuw ; Jedamzik:2010dq
@@
where $M_{\text{PBH},0}$ is the mean mass of the PBH distribution, considered as the mass of the PBH associated with the Gaussian peak (see eqns. (16) and (21)). Taking into account that for a successful BBN the temperature of the plasma should be $T_{\text{BBN}}>4\text{MeV}$ Hannestad:2004px and the upper limit on the energy scale of inflation, $H_{\text{inf}}/M_{\text{pl}}\lesssim 2.5\times 10^{-5}$ Planck:2018vyg ; Planck:2018jri , one obtains the following allowed range for the PBH masses, given by Domenech:2024wao ; Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2020ssp
@@ 3.1 Khlopov-Polnarev formalism
Figure 2: Fractional energy density of PBH for a) $k_{\text{peak}}=10^{-2}k_{\text{end}}$ and b) $k_{\text{peak}}=10^{-1}k_{\text{end}}$ as a function of the number of efolds $N$. Here, $k_{end}=5.7\times 10^{24}{\rm Mpc}^{-1}$ is the scale corresponding to the end of inflation, which exits the horizon approximately $N_{end}=60$ e-folds after the pivot scale. The vertical dashed line to the left corresponds to the point where the peak enters the horizon and the one to the left when the scale $k_{\text{peak}}/8$ enters the horizon. The horizontal dashed line corresponds to $\Omega_{\text{PBH}}(t)=1$.
@@ 3.2 Hawking evaporation
with $g_{H}$ being the spin-weighted degrees of freedom, which we assume to be $g_{H}\simeq 108$ (corresponding to standard model degrees of freedom) for $M_{\text{PBH}}\ll 10^{11}g$. In our case, we have an extended mass distribution, contrary to the standard monochromatic case. To simplify, we assume that the mean PBH mass corresponds to the mass of the PBH formed at the peak of the distribution, that is
@@ 3.4 Power spectrum of primordial black hole fluctuations
The fact that PBH are discrete objects introduces inhomogeneities, which can be quantified with the power spectrum of PBH density fluctuations $\delta_{\text{PBH},\bm{k}}$. For Poissonian statistics, it is given by Papanikolaou:2020qtd
@@
where $k_{\text{UV}}=a/\bar{r}$ is the ultraviolet cutoff of the power spectrum. For scales larger than $k_{\text{UV}}$, the PBH fluid can be treated as non-relativistic matter, but as $k$ approaches $k_{\text{UV}}$, the discrete nature of the PBH leads to shot-noise effects and the fluid picture is no longer valid Domenech:2023jve ; Domenech:2024wao . These energy density fluctuations correspond to isocurvature perturbations when the PBH form. In other words, PBH form in this case from the perturbations of the scalar field, and by conservation of energy, any missing part of the scalar field fluid that ends up into a PBH is compensated by PBH themselves, so that the fluctuations in both fluids are equal and opposite, i.e., $\delta_{\text{PBH},\bm{k}}=-\delta_{\phi,\bm{k}}$, which is what mainly characterizes isocurvature perturbations. Initially, these isocurvature PBH perturbations do not induce GWs, but as the PBH become the dominant component of the universe, the isocurvature PBH perturbations source curvature perturbations, which have a gravitational potential associated, and this last is responsible for inducing GWs. Note that these induced GWs should not be confused with the induced GWs from amplified perturbations during, for instance, an early matter-dominated period Domenech:2021ztg ; Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 ; draft-alpha-GWs . In this case, the GWs are sourced by the gravitational potential of the Poissonian fluctuations associated with the overproduction (dominance) of the PBH, instead of being sourced by the inflaton perturbations. In essence, the novelty of this approach resides in the fact that the isocurvature perturbations are sourced by the scalar field, contrary to the standard approach where the radiation fluid sources the isocurvature perturbations Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2024wao ; Domenech:2020ssp ; Papanikolaou:2022chm . Moreover, we do take into account the whole evolution of the PBH by considering their formation mechanism, instead of assuming an initial abundance of PBH.

### v2

@@ Abstract
In this work, we explore primordial black holes (PBH) formation scenario during the post-inflationary preheating stage dominated by the inflaton field. We consider, in particular, a model-independent parametrization of the Gaussian peak inflationary power spectrum that leads to amplified inflationary density fluctuations before the end of inflation. These modes can reenter the horizon during preheating and could experience instabilities that trigger the production of PBH. This is estimated with the Khlopov-Polnarev (KP) formalism that takes into account non-spherical effects. We derive an accurate analytical expression for the mass fraction under the KP formalism that fits well with the numerical evaluation. Particularly, we focus on ultra-light PBH of masses $M_{\text{PBH}}<10^{9}g$ and study their evolution and (possible) dominance after the decay of the inflation field into radiation and before the PBH evaporation via Hawking radiation. These considerations alter the previous estimates of induced gravitational waves (GWs) from PBH dominance and set new targets for detecting stochastic GW backgrounds with future detectors, provided that these achieve significantly enhanced experimental sensitivity, as current planned instruments do not yet possess sufficient sensitivity for detection.
@@ 1 Introduction
The inflationary phase constitutes the most important phase of the primordial Universe. It solves the conceptual challenges of the classical Big Bang Theory and provides a mechanism for generating the primordial density fluctuations observed in the cosmic microwave background (CMB). These fluctuations act as seeds for many phenomena, including primordial black holes (PBH) Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx ; Carr:1975qj and gravitational waves (GWs) Guzzetti:2016mkm ; Baumann:2009ds ; Maggiore:2007 ; Dodelson:2003 ; Gravitation:1973 ; Franciolini:2021 ; Domenech:2021ztg ; Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 ; Kohri:2018awv , in posterior eras. The specific nature of each era (matter or radiation-dominated) influences the dynamics of these phenomena, providing insights into early universe physics.
@@
Almost every inflationary model, especially single-field models, exhibits a phase known as preheating Kofman:1994rk ; Kofman:1997yn , just after the end of inflation. During this stage, the universe undergoes a transient matter-dominated period driven by the oscillations of the inflaton field at the bottom of the inflationary potential. PBH are usually considered to form from the collapse of amplified density perturbations generated during inflation as they enter the horizon. Conventionally, PBH formation has been studied in radiation-dominated eras, where the radiation pressure allows the definition of a collapse threshold for the perturbations Carr:2020gox ; Harada:2013epa ; Musco:2012au , and the mass fraction of PBH is often estimated using the Press-Schechter formalism Press:1973iz ; Harada:2013epa . In contrast, during this preheating stage, characterized by negligible pressure, the collapse dynamics differ significantly. Here, non-spherical effects become the dominant factor in stopping the collapse, a scenario well described by the Khlopov-Polnarev (KP) formalism Khlopov:1980mg ; Polnarev:1981 ; Khlopov:1982ef ; Polnarev:1985btg ; Khlopov:2008qy ; Harada:2016mhb . This last, typically ignored in the literature, was originally thought for matter-dominated scenarios. We choose to work under this formalism as we understand that it better describes the collapse dynamics of perturbations in matter-dominated scenarios. We have already explored this in previous works del-Corral:2023apl ; del-Corral:2025lcp , but here, we consider two new elements: the decay of the inflation field into radiation and the posterior evaporation of the PBH via Hawking radiation, which we describe in what follows. Further, we will apply these considerations to an extended distribution of perturbations, contrary to the standard monochromatic assumption.
@@
Since PBH form from the collapse of amplified perturbations, a central element in its formation is the appearance of a pronounced peak in the inflationary power spectrum. Various inflationary scenarios can lead to such an amplification of curvature perturbations. For instance, in single-field models, mechanisms such as an inflection point Iacconi:2021ltm ; Dalianis:2018frf ; Garcia-Bellido:2017mdw ; Hertzberg:2017dkh or a step-like feature in the potential Inomata:2021tpx ; Dalianis:2023pur ; Dalianis:2021iig ; Cai:2021zsp ; Papanikolaou:2022did can generate the necessary peak. For multi-field, this peak can be produced through non-canonical kinetic terms Aldabergenov:2022rfc ; Braglia:2020eai ; Pi:2021dft ; Wang:2024vfv , the interplay of multiple axion fields Zhou:2020kkf ; Mavromatos:2022yql , waterfall trajectories in hybrid models Afzal:2024xci ; Spanos:2021hpk ; Braglia:2022phb ; Clesse:2015wea ; Tada:2023fvd or rapid turns in the inflationary trajectory Fumagalli:2020adf ; Palma:2020ejf ; Anguelova:2020nzl . For a comprehensive review of these mechanisms, see Stamou:2024lqf . Although the parametrization of the power spectrum via a Gaussian peak is introduced in an ad hoc manner in our study, adjusting its height and position effectively encapsulates the essential physics of almost every inflationary model mentioned earlier. This is the approach we will follow, remaining agnostic about the precise physical mechanism of the amplification of perturbations during inflation.
@@
Interest in PBH has grown recently due to their potential to answer several questions in cosmology Stojkovic:2004hz ; Stojkovic:2005zh . In particular, they have been proposed as candidates for dark matter, as generators of structure in the universe, or even as seeds for the formation of supermassive black holes in the center of galactic nuclei. See Villanueva-Domingo:2021spv ; Carr:2020gox for a review. However, although observational data tightly constrain their abundance Carr:2020gox ; Cai:2021zsp ; Markov:1984xd ; Coleman:1991ku ; Carr:1994ar ; Zeldovich ; Acharya:2020jbv ; Carr:2009jm ; Josan:2009qn ; Carr:2020xqk , for small masses ($\lesssim 10^{10}$g), these constraints rely on the nature of Dark Matter, which is currently unknown Allahverdi:2020bys . For this reason, the constraints can be relaxed to the point that PBH come to dominate the energy density of the universe, a period called PBH-dominated, which behaves as pressureless matter. If this occurs, their inherent Poissonian fluctuations in density can source a stochastic background of gravitational waves, providing a unique observational window into the early universe. The frequencies associated with this background of GWs fall into the very-high frequency range, as shown in previous works in the context of PBH in grand unified theories and supergravity Anantua:2008am ; Zagorac:2019ekv ; del-Corral:2025fzz . In this range, the current and planned detectors still lack a sufficiently high sensitivity, further motivating their refinement.
@@
To eventually recover the standard radiation-dominated era necessary for successful Big Bang Nucleosynthesis (BBN), we consider the inflaton decay into Standard Model particles, thereby reheating the universe. This is achieved by considering the scalar field’s decay into radiation, with decay rate $\Gamma_{\phi}$. However, if PBH dominates before that, then the universe is reheated through Hawking evaporation Hawking:1974rv ; Martin:2019nuw ; Gondolo:2020uqv ; Dong:2015yjs ; Holst:2024ubt , and the reheating temperature corresponds to the evaporation temperature of the PBH. Refs. Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2024wao ; Domenech:2020ssp ; Papanikolaou:2022chm ; Papanikolaou:2024kjb ; He:2024luf ; Inomata:2019ivs considered that the PBH-dominated phase occurs due to the interplay between the different energy densities of PBH (matter), and a background fluid with an arbitrary equation of state $w>0$, ignoring the specific formation mechanism of the PBH. We, however, consider that PBH are produced during an early matter-dominated phase, under the KP formalism, and that the PBH-dominated phase occurs due to the decay of the scalar field into radiation, which modifies the actual constraints on the abundance of PBH. Then, we analyze the production of induced GWs by the Poissonian fluctuations of the PBH fluid. To our knowledge, this is the first time it has been studied in the literature.
@@ 2 Inflation and preheating
where $\mathcal{A}_{s}=2.1\times 10^{-9}$ is the amplitude of the power spectrum at the pivot scale $k_{0}=0.05\text{Mpc}^{-1}$, and $n_{s}\simeq 0.965$ is the spectral index. Motivated by other works Iacconi:2021ltm ; Dalianis:2018frf ; Stamou:2024lqf ; Garcia-Bellido:2017mdw ; Hertzberg:2017dkh ; Inomata:2021tpx ; Dalianis:2023pur ; Dalianis:2021iig ; Cai:2021zsp ; Aldabergenov:2022rfc ; Braglia:2020eai ; Pi:2021dft ; Wang:2024vfv ; Zhou:2020kkf ; Mavromatos:2022yql ; Afzal:2024xci ; Spanos:2021hpk ; Braglia:2022phb ; Clesse:2015wea ; Tada:2023fvd ; Papanikolaou:2022did ; Fumagalli:2020adf ; Palma:2020ejf ; Anguelova:2020nzl (see the introduction), we introduce a Gaussian peak in the power spectrum, and choose to parametrize it as a log-normal-like distribution, given by
@@
where $\mathcal{A}_{\text{peak}}$ is the amplitude at the peak scale $k_{\text{peak}}$ and the standard deviation of the Gaussian is chosen to be $\sigma=0.2$ to reduce the number of free parameters. Eqn. (2) is useful for this work as it avoids the tails of the Gaussian peak (due to the $\log_{10}$ term), and since it constitutes a two-parameter ($\mathcal{A}_{\text{peak}}$ and $k_{\text{peak}}$) model-independent parametrization that can reflect the physics of the models described in Iacconi:2021ltm ; Dalianis:2018frf ; Garcia-Bellido:2017mdw ; Hertzberg:2017dkh ; Inomata:2021tpx ; Dalianis:2023pur ; Dalianis:2021iig ; Cai:2021zsp ; Aldabergenov:2022rfc ; Braglia:2020eai ; Pi:2021dft ; Wang:2024vfv ; Zhou:2020kkf ; Mavromatos:2022yql ; Afzal:2024xci ; Spanos:2021hpk ; Braglia:2022phb ; Clesse:2015wea ; Tada:2023fvd ; Stamou:2024lqf ; Papanikolaou:2022did ; Fumagalli:2020adf ; Palma:2020ejf ; Anguelova:2020nzl . The total power spectrum is the sum of the contributions from (1) and (2), that is
@@
During this phase, the perturbations that enter the horizon collapse into PBH, as shown in the next section. However, the quantity we use to compute the abundance of PBH, i.e., the mass fraction $\beta(k)$, is not the power spectrum of curvature perturbations, but the variance of the density perturbations, $\sigma_{k}$. For modes that enter the horizon after inflation, the curvature perturbations are related to the density perturbations $\delta_{\phi,\bm{k}}$ as follows  [^1] In this work, since our focus is only on the scales that are amplified during inflation, we do not consider the subhorizon modes that never exit the horizon during inflation. Although they could contribute to PBH formation, as shown in del-Corral:2023apl ; del-Corral:2025lcp , their contribution is only toward very small scales in contrast to what we explore in this work.Martin:2019nuw ; Jedamzik:2010dq
@@
with $g_{H}$ being the spin-weighted degrees of freedom and $M_{\text{PBH},0}$ the mean mass of the PBH distribution, considered as the mass of the PBH associated with the Gaussian peak (see eqns. (16) and (21)). Taking into account that for a successful BBN the temperature of the plasma should be $T_{\text{BBN}}>4\text{MeV}$ Hannestad:2004px ; Kawasaki:2000en ; Hasegawa:2019jsa and the upper limit on the energy scale of inflation, $H_{\text{inf}}/M_{\text{pl}}\lesssim 2.5\times 10^{-5}$ Planck:2018vyg ; Planck:2018jri , one obtains the following allowed range for the PBH masses, given by Domenech:2024wao ; Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2020ssp
@@ 3.1 Khlopov-Polnarev formalism
Figure 2: Fractional energy density of PBH for a) $k_{\text{peak}}=10^{-2}k_{\text{end}}$ and b) $k_{\text{peak}}=10^{-1}k_{\text{end}}$ as a function of the number of efolds $N$. Here, $k_{end}=5.7\times 10^{24}{\rm Mpc}^{-1}$ is the scale corresponding to the end of inflation, which exits the horizon approximately $N_{end}=60$ e-folds after the pivot scale. The vertical dashed line to the left corresponds to the point where the peak enters the horizon, and the one to the right when the scale $k_{\text{peak}}/8$ enters the horizon. The horizontal dashed line corresponds to $\Omega_{\text{PBH}}(t)=1$.
@@ 3.2 Hawking evaporation
where we assume $g_{H}\simeq 108$ (corresponding to standard model degrees of freedom) for $M_{\text{PBH}}\ll 10^{11}g$. In our case, we have an extended mass distribution, contrary to the standard monochromatic case. To simplify, we assume that the mean PBH mass corresponds to the mass of the PBH formed at the peak of the distribution, that is
@@ 3.4 Power spectrum of primordial black hole fluctuations
The fact that PBH are discrete objects introduces inhomogeneities, which can be quantified with the power spectrum of PBH density fluctuations $\delta_{\text{PBH},\bm{k}}$. For Poissonian statistics, it is given by [^3] In this paper, we are not considering possible deviation from Poisson statistics by non-Gaussianities that could further lead to primordial clustering Papanikolaou:2024kjb ; He:2024luf because our consideration of the KP collapse mechanism is restricted to overdensities that are small enough. Papanikolaou:2020qtd
@@
where $k_{\text{UV}}=a/\bar{r}$ is the ultraviolet cutoff of the power spectrum. For scales larger than $k_{\text{UV}}$, the PBH fluid can be treated as non-relativistic matter, but as $k$ approaches $k_{\text{UV}}$, the discrete nature of the PBH leads to shot-noise effects and the fluid picture is no longer valid Domenech:2023jve ; Domenech:2024wao . These energy density fluctuations correspond to isocurvature perturbations when the PBH form. In other words, PBH form in this case from the perturbations of the scalar field, and by conservation of energy, any missing part of the scalar field fluid that ends up into a PBH is compensated by PBH themselves, so that the fluctuations in both fluids are equal and opposite, i.e., $\delta_{\text{PBH},\bm{k}}=-\delta_{\phi,\bm{k}}$, which is what mainly characterizes isocurvature perturbations. Initially, these isocurvature PBH perturbations do not induce GWs, but as the PBH become the dominant component of the universe, the isocurvature PBH perturbations source curvature perturbations, which have a gravitational potential associated, and this last is responsible for inducing GWs. Note that these induced GWs should not be confused with the induced GWs from amplified perturbations during, for instance, an early matter-dominated period Domenech:2021ztg ; Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 ; del-Corral:2025fzz . In this case, the GWs are sourced by the gravitational potential of the Poissonian fluctuations associated with the overproduction (dominance) of the PBH, instead of being sourced by the inflaton perturbations. In essence, the novelty of this approach resides in the fact that the isocurvature perturbations are sourced by the scalar field, contrary to the standard approach where the radiation fluid sources the isocurvature perturbations Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2024wao ; Domenech:2020ssp ; Papanikolaou:2022chm . Moreover, we do take into account the whole evolution of the PBH by considering their formation mechanism, instead of assuming an initial abundance of PBH.